Find Particular solution of d²y/dx² -y=sinh2x
1
Expert's answer
2020-12-06T19:54:14-0500

Let us find the solution of the equation "\frac{d²y}{dx²} -y=\sinh2x". Note that "\sinh 2x =\frac{e^{2x}-e^{-2x}}{2}" .


The characteristic equation "k^2-1=0" of the equation "\frac{d²y}{dx²} -y=0" has the solutions "k=1"and "k=-1."


The solution of the equation "\frac{d²y}{dx²} -y=\frac{e^{2x}-e^{-2x}}{2}" is "y=C_1e^x+C_2e^{-x+}y_p", where "y_p" is a particular solution of "\frac{d²y}{dx²} -y=\frac{e^{2x}-e^{-2x}}{2}" .


"y_p=Ae^{2x}+Be^{-2x}"


"y_p'=2Ae^{2x}-2Be^{-2x}"


"y_p''=4Ae^{2x}+4Be^{-2x}"


"4Ae^{2x}+4Be^{-2x}-(Ae^{2x}+Be^{-2x})=\frac{e^{2x}-e^{-2x}}{2}"


"3Ae^{2x}+3Be^{-2x}=\frac{1}{2}e^{2x}-\frac{1}{2}e^{-2x}"


"3A=\frac{1}{2}" and "3B=-\frac{1}{2}"


"A=\frac{1}{6}" and "B=-\frac{1}{6}"


Therefore, the general solution is the following:


"y=C_1e^x+C_2e^{-x}+\frac{1}{6}e^{2x}-\frac{1}{6}e^{-2x}=C_1e^x+C_2e^{-x}+\frac{1}{3}\sinh 2x"


The particular solution is "y_p=\frac{1}{3}\sinh 2x" (for "C_1=C_2=0").